If C = [2.5 cm, 80°], i.e., the magnitude and direction of C are 2.5 cm and 80°, D = [3.5 cm,

120°], and E = D – 2C, then what is the direction of E (to the nearest degree)?

 


1
Expert's answer
2021-10-06T10:05:25-0400
"\vec c=(2.5\cos 80\degree, 2.5\sin 80\degree)"

"\vec d=(3.5\cos 120\degree, 3.5\sin 120\degree)"

"\vec e=\vec d-2\vec c=(3.5\cos 120\degree, 3.5\sin 120\degree)"

"-2(2.5\cos 80\degree, 2.5\sin 80\degree)"

"=(3.5\cos 120\degree-5\cos 80\degree, 3.5\sin 120\degree-5\sin 80\degree)"

"\tan \theta=\dfrac{3.5\sin 120\degree-5\sin 80\degree}{3.5\cos 120\degree-5\cos 80\degree}"

"\theta=\tan^{-1}(\dfrac{3.5\sin 120\degree-5\sin 80\degree}{3.5\cos 120\degree-5\cos 80\degree})"

"\approx36\degree"